The lengths of rods from a certain factory are normally distributed with mean xcm and standard deviation 6cm. It is known that 4.78% of the rods have lengths greater than 82cm. Find to the nearest integer,
(i) the mean x.
(ii) the range, s (symmetrical about the mean) within which 75% of the lengths of the rods lie.
(iii) the probability that a rod chosen at random has a length between 62cm and 72cm.
1
Expert's answer
2021-01-19T04:15:05-0500

(i) Let "Y=" the length of rod: "Y\sim N(x, \sigma^2)." Then "Z=\dfrac{Y-x}{\sigma}\sim N(0, 1)"

Given "\sigma=6\ cm."


"P(Y>82)=1-P(Y\leq82)"

"=1-P(Z\leq\dfrac{82-x}{6})=0.0478"

"P(Z\leq\dfrac{82-x}{6})=0.9522"

"\dfrac{82-x}{6}\approx1.66657"

"x\approx72"

"mean=x=72"


(ii)

"P(x-\dfrac{s}{2}<Y<x+\dfrac{s}{2})"

"=P(Y<x+\dfrac{s}{2})-P(Y\leq x-\dfrac{s}{2})="

"=P(Z<\dfrac{\dfrac{s}{2}}{6})-P(Z\leq\dfrac{-\dfrac{s}{2}}{6})=0.75"

"P(Z\leq-\dfrac{s}{12})=\dfrac{1-0.75}{2}"

"-\dfrac{s}{12}\approx-1.15035"

"s\approx13.0842"

(iii)


"P(62<Y<72)=P(Y<72)-P(Y\leq62)"

"=P(Z<\dfrac{72-72}{6})-P(Z\leq\dfrac{62-72}{6})"

"=P(Z<0)-P(Z\leq-\dfrac{5}{3})\approx0.5-0.04779"

"=0.45221"

The probability that a rod chosen at random has a length between 62cm and 72cm is 0.45221.